March BG Com Night 11, UT 2009 March 17

Observers: Dave Thompson (@LBT)
Telescope Operator: D Gonzalez Huerta
Lucifer support: N Ageorges (@LBT)
SW Support: Paul Grenz
Telescope Support: D Officer, J Little

Summary:

  • PSF/Collimation problems, see IT#2072
  • Deep NB exposure w/LUCIFER (5 minutes each position)
  • Took ~72 point dataset on pointing model (last night's was not good)
  • Collimation stayed pretty good through the whole 2 hours
  • Empirical correction to oacontrol transformation failed.

Details:

Dave loaded new pointing model based on pointing data from last night.

IE/CA adjustment

UT 2:05:29: On a couple of presets, Lucifer receives the following error message: oss.side[0].tertiary_mirror:[adjustCurrFocalStation] Not at a focal station. IssueTrak #2071.

MST 20:02 Preset would not return because collimation would not converge with lots of PSF error messages. Paul IssueTrak'd as #2072

Running AGW offset testing

MST 20:54 Lucifer team noticed that we hit -Y limit on primary mirror. Added +1 mm to Yglobal for both M1 and M2, although took a couple tries to get it right. Lost ~10min

MST 21:25 Lots of error messages in log "primary mirror collimation position error" caused by cranky hardpoints. This is a common problem, especially in cold weather. Fixed by a forced (manual) re-start of collimation cycle. No observing time lost.

MST 22:13 Starting check of new pointing model. Note that the rms of the fit was poor - 8 arcsec - but this could be because the data were taken out of equilibrium.

MST 22:21 pointto WT10_350 PA=0 gs=0 Tel="ACQUIRE" PAM-
Correct pointing: IE changed from -7.91 to -17.50, CA from 11.69 to 10.50 at EL~72
MST 22:24 pointto WT10_403 PA=0 gs=0 Tel="ACQUIRE", star off edge of acq image, EL~20
MST 22:26 pointto ACT0402 PA=0 gs=0 Tel="ACQUIRE", star near bottom of image about 10 arcsec off, EL~36
MST 22:28 pointto WT10_359 PA=0 gs=0 Tel="ACQUIRE", off ~7 arcsec, correcting pointing again here at EL~60.
IE now -24, CA now 13.0

Pointing sufficiently bad that I will take another set of data, but the seening is quite good (0.4-0.5 on guider) so LUCIFER is taking some long integrations (NB data, OH filter between lines in J). Use these data to extract guiding and WFS rms, overheads on dithering, etc.

luci_20090316_ptg_NNNN
0002-0005 are rotator center at PA=0, 180, 90, 270 degrees
pointing log file = /home/telescope/20090317.ptlog
MST 02:27 = UT 09:27 at start of run
StarID NNNN StarID NNNN StarID NNNN
WT10_354 0006 WT10_394 0030 ACT0287 0054
WT10_405 0007 ACT0407 0031 ACT0326 0055
WT10_404 0008 WT10_411 0032 WT10_317 0056
WT10_363 0009 ACT0341 0033 WT10_347 0057
ACT0288 0010 ACT0382 0034 WT10_334 0058
WT10_186 0011 ACT0403 0035 ACT0358 0059
ACT0260 0012 WT10_409 0036 WT10_343 0060
WT10_088 0013 ACT0365 0037 ACT0315 0061
WT10_175 0014 WT10_381 0038 ACT0268 0062
ACT0264 0015 WT10_350 0039 WT10_215 0063
ACT0215 0016 WT10_363 0040 WT10_330 0064
ACT0317 0017 ACT0294 0041 WT10_162 0065
WT10_330 0018 WT10_205 0042 WT10_356 0066
WT10_300 0019 WT10_169 0043 ACT0405 0067
WT10_249 0020 WT10_171 0044 WT10_399 0068
WT10_296 0021 ACT0190 0045 WT10_290 0069
ACT0192 0022 ACT0245 0046 WT10_171 0070
WT10_077 0023 ACT0192 0047 WT10_245 0071
WT10_137 0024 ACT0141 0048 WT10_090 0072
WT10_279 0025 ACT0380 0049 ACT0160 0073
ACT0241 0026 ACT0109 0050 ACT0283 0074
WT10_220 0027 WT10_196 0051 WT10_217 0075
WT10_290 0028 WT10_249 0052 WT10_319 0076
ACT0326 0029 ACT0283 0053 ACT0341 0077

Collimation stayed remarkably good throughout the whole pointing run. We took advantage of the stability to collect a good set of data.

MST 0446 Preset to BS9158, gs=0 to collimate. WFS rms = 3068.56, 1022.69, 404.45...
See acquire000024 for first image after 2.25 hour pointing run.

MST04:53 Preset to BS9153 gs=15 to get PA for next test (156.22)
Pointing correction: IE from -23.88 to -30.00, CA from 11.32 to 20.00. Tweaked the focus +0.1 mm Zglobal.

Test 0 (default config): luci_20090316_agw2_NNNN
GS NNNN
15 0001
6  

Trouble finding off-axis stars (they are on acq image but do not make it to the guide box).

Tue Mar 17 05:23:07.438 2009 Preset alert for left LUCIFER mode ACQUIRE: PCS could not start tracking because ERROR: MCSPU is not connected.
DGH killed and restarted PCS.

MST 05:31 try again:
GS NNNN
15 0002
16 0003
6 0004

both 6 and 16 showed up right at the edge of the guide box. Behavior as before, on-axis source is displaced between images.

Change psi_gpoff from 2.8353 to -2.8353 and recheck:
GS NNNN
15 0005

failed to acquire 16 (at edge of guide box) and also failed to acquire 6, same reason. The oacontrol.conf file was restored to how it was prior to this test.

Sky too bright to continue. Turning over to LUCIFER for twilight flats.

-- DavidThompson - 16 Mar 2009
Topic revision: r9 - 27 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback